# An Ecosystem for Ontology Interoperability [Vision]

### v1

@@
Abstract.
Ontology interoperability is one of the complicated issues that restricts the use of ontologies in knowledge graphs (KGs). Different ontologies with conflicting and overlapping concepts make it difficult to design, develop, and deploy an interoperable ontology for downstream tasks. We propose an ecosystem for ontology interoperability. The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability in real-world applications. A case study in the building domain validates the usefulness of the proposed ecosystem.
PVLDB Reference Format:
PVLDB, 19(1): XXX-XXX, 2025.
doi:$X$.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 19, No. 1 ISSN 2150-8097.
doi:$X$.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/qzc438-research/ontology-interoperability.
1. Introduction
For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability between KGs. While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs. For example, it can be challenging to understand a KG snippet (Sydney, visit, Sydney) alone, but with a logic-based expression (Person, TravelAction, Place) derived from the schema.org vocabulary (Schema.org, 2025), one can infer that the subject “Sydney” refers to a visitor, while the object “Sydney” refers to the place being visited. The triple (Sydney, visit, Sydney) describes an event in which the American actress Sydney Sweeney visited the city of Sydney (Mitchell, 2023).
@@ 1. Introduction
In this paper, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs). We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
2. Related Work
2.1. Ontology Design Patterns
The goal of ontology design patterns (ODPs) is to explore common and shared ontology fragments between different ontologies (Gangemi and Presutti, 2009). Figure 1 illustrates that ODPs use the star-based matching schema. Once the ODP is determined, each new ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of the ontology.
Figure 1. Ontology design patterns (ODPs).
[…]
Ontology design patterns (ODPs).
@@ 2.1. Ontology Design Patterns
2.2. Ontology Matching and Versioning
Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies (Euzenat and Shvaiko, 2013) or different versions of the same ontology (Klein and Fensel, 2001). Figure 2 shows that OM&OV conform to the pairwise matching schema. This means that each mapping is a pairwise process to find the possible correspondences between two ontologies.
Figure 2. Ontology matching and versioning (OM&OV).
[…]
Ontology matching and versioning (OM&OV).
@@ 2.2. Ontology Matching and Versioning
2.3. Ontology-Compliant Knowledge Graphs
Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment based on the given KG(s) (Qiang, 2023). Figure 3 shows that the OCKGs comply with the data-driven schema. While ontologies comply with KG(s), their reference mappings can be derived from similar concepts pointing to the same data point.
Figure 3. Ontology-compliant knowledge graphs (OCKGs).
[…]
Ontology-compliant knowledge graphs (OCKGs).
@@ 2.3. Ontology-Compliant Knowledge Graphs
3. Our Vision: An Ecosystem
Figure 4 illustrates an overview of the ecosystem for ontology interoperability. Similarly to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve better interoperability.
Figure 4. An overview of the ecosystem.
[…]
An overview of the ecosystem.
@@ 3. Our Vision: An Ecosystem
4. A Case Study in the Building Domain
Driven by the emergence of efficient digital transformation for Industry 4.0, numerous building ontologies have been developed over the past two decades. While building concepts are complex and evolve rapidly, ontology interoperability has become a prominent problem in the building domain (Qiang et al., 2023). We use a running example to illustrate the usefulness of our proposed ecosystem for ontology interoperability. The running example is described as follows: A sensor is installed in the room to monitor the room temperature. The ontologies used in this case are required to capture several main concepts, including sensor, room, temperature, and sensor observation.
4.1. Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is recommended to build an ontology based on ODPs. Several ODPs can be used in this case (shown in Figure 5).
@@ 4.1. Design Phase
    Observation Pattern (used in SOSA/SSN (Haller et al., 2017)). Sensor observation is defined as a type of observation. The observation is made by the sensor, which has a feature of interest in the room and the observed property temperature.
@@
    Measurement Pattern (used in the early version of SAREF (Daniele et al., 2015)). Sensor observation is defined as a measurement. The measurement is made by the sensor, which is a measure of the room related to the property temperature.
@@
    Event Pattern (used in DUL (Presutti and Gangemi, 2016)). Sensor observation is defined as an event. The event is associated with the sensor, room, and temperature.
@@
    Activity Pattern (used in PROV-O (Lebo et al., 2013)). Sensor observation is defined as an activity. The activity is associated with the sensor that uses the room and generates a temperature measurement.
Figure 5. Modelling using different ODPs.
[…]
Modelling using different ODPs.
@@
4.2. Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and SSN. As illustrated in Figure 6, we can see that SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to find aligned concepts between two different ontologies. If we consider SOSA to be an evolved version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV tasks can be unified with minor modifications. OM&OV tasks can be classified into two main categories.
Figure 6. Aligning SOSA and SSN.
[…]
Aligning SOSA and SSN.
  1. (1)
@@ 4.2. Develop Phase
  2. (2)
    Finding non-trivial alignments considers syntactic, lexical, and semantic matching of the entity. For example, the mappings between the classes sosa:ObservableProperty and ssn:Property, and the properties sosa:isHostedBy and ssn:onPlatform. The alignment should also exclude non-existent and counter-intuitive mappings. For example, ssn:hasProperty should not have a direct SOSA equivalent mapping.
4.3. Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it remains challenging to determine which ontology to use in the real world. Therefore, we need to apply these two ontologies in the KG(s) and construct the OCKGs for evaluation. Figure 7 shows the OCKGs based on SOSA and SSN. We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the KG instance.
  1. (1)
    The stimulus “Exceeds 25 Degrees”. The sensor only makes an observation when the temperature exceeds 25 degrees.
  2. (2)
@@ 4.3. Deploy Phase
To model these two new concepts, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA. We also make several design choices between aligned mappings.
  1. (1)
@@
  3. (3)
    We use sosa:hasFeatureOfInterest to replace ssn:FeatureOfInterest. Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we distinguish these two concepts to avoid confusion.
Figure 7. Build OCKGs using SOSA and SSN.
[…]
Building OCKGs using SOSA and SSN.
Figure 8 shows the evolved ontology and the OCKG built on the evolved ontology. Apparently, the evolved ontology is more suitable for the provided KG instance. In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times. It is also possible to choose other alternative ODPs. For example, the class “BuildingSpace” in the Measurement Pattern is more suitable to describe a room, and the property “hasLocation/isPartOf” in the Event Pattern is more appropriate to describe the link between the sensor and the room.
Figure 8. Evolved ontology and new OCKG.
[…]
Evolved ontology and new OCKG.
4.4. Implications
Well-defined ODPs can be used to create reusable ontology fragments, thereby reducing the complexity of OM&OV tasks. As a complement to ODPs, OM&OV can align the different concepts that are not captured by ODPs. On the other hand, while OCKGs are useful tools for validating ontology interoperability and evolving ODPs, the construction of OCKGs is heavily based on the precise mapping produced by OM&OV systems. OM&OV is a critical middleware between ODPs and OCKGs.
5. Discussion & Conclusion
@@
Acknowledgements.

### v2

@@
Abstract
Ontology interoperability is one of the complicated issues that restricts the use of ontologies in knowledge graphs (KGs). Different ontologies with conflicting and overlapping concepts make it difficult to design, develop, and deploy an interoperable ontology for downstream tasks. We propose an ecosystem for ontology interoperability. The ecosystem employs three state-of-the-art semantic techniques in different phases of the ontology engineering life cycle: ontology design patterns (ODPs) in the design phase, ontology matching and versioning (OM&OV) in the develop phase, and ontology-compliant knowledge graphs (OCKGs) in the deploy phase, to achieve better ontology interoperability in real-world applications. A case study of sensor observation in the building domain validates the usefulness of the proposed ecosystem.
Index Terms:
ontology, data interoperability
I Introduction
For large-scale knowledge graphs (KGs), the use of ontologies has become the general practice for enabling semantic interoperability between KGs. While building KGs does not always require an ontology, ontologies can provide extensive information to understand KGs. For example, it can be challenging to understand a KG snippet (Sydney, travelsTo, Sydney) alone, but with a logic-based expression (Person, travelsTo, Place), one can infer that the subject “Sydney” refers to a visitor, while the object “Sydney” refers to the place being visited. The triple (Sydney, travelsTo, Sydney) describes an event in which the American actress Sydney Sweeney travels to the city of Sydney [1].
@@ I Introduction
In this paper, we review three state-of-the-art semantic web techniques: ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs). We compare the differences and connections between these three techniques and explore their roles in ontology engineering to improve interoperability. We propose an ecosystem for ontology interoperability and use a case study of sensor observation in the building domain to showcase how these three techniques can complement each other’s strengths and mitigate their weaknesses within the ontology engineering life cycle.
II Related Work
II-A Ontology Design Patterns
The goal of ontology design patterns (ODPs) is to explore common and shared ontology fragments between different ontologies [2]. Fig. 1 illustrates that ODPs use the star-based matching schema. Once the ODP is determined, each subsequent ontology will conform to this ODP, and the ODP serves as a high-level conceptual model to guide the design of these new ontologies.
Figure 1: Ontology design patterns (ODPs).
@@ II-A Ontology Design Patterns
II-B Ontology Matching and Versioning
Ontology matching and versioning (OM&OV) aim to find reference mappings between different ontologies [3] or different versions of the same ontology [4]. Fig. 2 shows that OM&OV conform to the pairwise matching schema. This means that each mapping is a pairwise process to find the possible correspondences between two ontologies.
Figure 2: Ontology matching and versioning (OM&OV).
@@ II-B Ontology Matching and Versioning
II-C Ontology-Compliant Knowledge Graphs
Ontology-compliant knowledge graphs (OCKGs) use a pay-as-you-go model to find the most cost-effective ontology fragment based on the given KG(s) [5]. Fig. 3 shows that the OCKGs comply with the star-based matching schema. While ontologies comply with KG(s), their reference mappings can be derived from the concepts pointing to the same data instance.
Figure 3: Ontology-compliant knowledge graphs (OCKGs).
@@ II-C Ontology-Compliant Knowledge Graphs
III An Ecosystem
Fig. 4 illustrates an overview of the ecosystem for ontology interoperability. Similarly to software engineering, ontology engineering involves three main phases: design, develop, and deploy. We employ ontology design patterns (ODPs), ontology matching and versioning (OM&OV), and ontology-compliant knowledge graphs (OCKGs) in different phases of the ontology engineering life cycle to achieve better interoperability.
Figure 4: An overview of the ecosystem.
@@ III An Ecosystem
IV A Case Study in the Building Domain
Driven by the emergence of efficient digital transformation for Industry 4.0, numerous building ontologies have been developed over the past two decades. While building concepts are complex and evolve rapidly, ontology interoperability has become a prominent problem in the building domain [6]. We use a running example to illustrate the usefulness of our proposed ecosystem for ontology interoperability. The running example is described as follows: A sensor is installed in the room to monitor the room temperature. The ontologies used in this case are required to capture several main concepts, including sensor, room, temperature, and sensor observation.
IV-A Design Phase
Ontologies can be built from scratch, but such ontologies are likely to have more customised concepts and limited interoperability. It is recommended to build an ontology based on ODPs. Several ODPs can be used in this case, as illustrated in Fig. 5.
@@ IV-A Design Phase
    Observation Pattern (used in SOSA [7] and SSN [8]). Sensor observation is defined as a type of observation. The observation is made by the sensor, which has a feature of interest in the room and the observed property temperature.
@@
    Measurement Pattern (used in the early version of SAREF [9]). Sensor observation is defined as a measurement. The measurement is made by the sensor, which is a measure of the room related to the property temperature.
@@
    Event Pattern (used in DUL [10]). Sensor observation is defined as an event. The event is associated with the sensor, room, and temperature.
@@
    Activity Pattern (used in PROV-O [11]). Sensor observation is defined as an activity. The activity is associated with the sensor that uses the room and generates a temperature measurement.
@@
Figure 5: Modelling using different ODPs.
IV-B Develop Phase
We assume that the Observation Pattern is chosen in the design phase and implemented in two ontologies, SOSA and SSN. As illustrated in Fig. 6, we can see that SOSA and SSN are very similar because they are built from the same ODPs. However, they are not exactly the same. The OM task is to find aligned concepts between two different ontologies. If we consider SOSA to be an updated version of SSN, then this OM task is also an OV task. We use OM&OV here because OM and OV can be unified with minor modifications [12]. OM&OV tasks can be classified into two main categories.
Figure 6: Aligning SOSA and SSN.
  1. 1.
@@ IV-B Develop Phase
  2. 2.
    Finding non-trivial alignments considers syntactic, lexical, and semantic matching of the entity. For example, the mappings between the classes sosa:ObservableProperty and ssn:Property, and the properties sosa:isHostedBy and ssn:onPlatform. The alignment should also exclude non-existent and counter-intuitive mappings. For example, ssn:hasProperty should not have a direct equivalent mapping in SOSA.
IV-C Deploy Phase
Now we have developed SOSA and SSN, along with their reference alignment. However, it remains challenging to determine which ontology to use in the real world. Therefore, we need to apply these two ontologies in the KG(s) and construct the OCKGs for evaluation. Fig. 7 shows the OCKGs based on SOSA and SSN. We can see that both ontologies are able to capture the general concepts of this sensor observation; however, they fail to capture two specific concepts in the KG instance.
  1. 1.
    The stimulus of the sensor “Exceeds 25 Degrees”. The sensor only makes an observation when the temperature exceeds 25 degrees.
  2. 2.
@@ IV-C Deploy Phase
To model these two new concepts, we need to extend the ODP with the Stimulus-Sensor-Observation Pattern from SSN and the Sampling Pattern from SOSA. We also make several design choices between aligned mappings (retrieved from [13, 14, 15]).
  1. 1.
@@
  3. 3.
    We use sosa:hasFeatureOfInterest to replace ssn:FeatureOfInterest. Although there is no restriction for ssn:FeatureOfInterest to be used both as a class and as a property, we distinguish them to avoid confusion.
Figure 7: Build OCKGs using SOSA and SSN.
Fig. 8 shows the updated ontology and the OCKG built on the updated ontology. Apparently, the updated ontology is more suitable for the provided KG instance. In the real world, an ontology can be deployed in multiple KGs, and the procedure described above can be repeated multiple times. It is also possible to choose other alternative ODPs. For example, the class “BuildingSpace” in the Measurement Pattern is more suitable to describe a room, and the property “hasLocation/isPartOf” in the Event Pattern is more appropriate to describe the link between the sensor and the room.
Figure 8: Updated ontology and new OCKG.
IV-D Implications
Well-defined ODPs can be used to create reusable ontology fragments, thereby reducing the complexity of OM&OV tasks. As a complement to ODPs, OM&OV can align the concepts that are not captured by ODPs. On the other hand, while OCKGs are useful tools for validating ontology interoperability and evolving ODPs, the construction of OCKGs is heavily based on the precise mapping produced by OM&OV systems. OM&OV is a critical middleware between ODPs and OCKGs.
V Discussion & Conclusion
@@
Supplemental Material Statement
The source code, data, and/or other artifacts used in this study have been made available at https://github.com/qzc438-research/ontology-interoperability.
Acknowledgment